OpenStudy (anonymous):

simplify given that all angles are between 0 and pi Cos(2arccos sqrt2/2)

OpenStudy (mertsj):

\[\cos^{-1} \frac{\sqrt{2}}{2}=\frac{\pi}{4}, \frac{7\pi}{4}\]

OpenStudy (mertsj):

When you multiply those angles by 2 you get: \[\frac{\pi}{2},\frac{7\pi}{2}\]

OpenStudy (mertsj):

\[\cos \frac{\pi}{2}=0; \cos \frac{7\pi}{2}=0\]